# Attractors for Singular-Degenerate Porous Medium Type Equations Arising in Models for Biofilm Growth

### v1

@@ 2.2 Long-time Behaviour of Solutions
$\mathcal{X}:=\overline{\{u\in L^{1}(\Omega):\phi(u)\in L^{\infty}(\Omega)\}},$
where the closure is taken in $L^{1}(\Omega)$. Then, by Theorem 2.1, problem (2.1) generates a semigroup $\{\mathcal{S}(t)\}_{t\geq 0}$ in $\mathcal{X}$ that is Lipschitz continuous in $L^{1}(\Omega)$. More specifically, $\mathcal{S}(t)u_{0}=u(t)$, where $u\in C([0,T];L^{1}(\Omega))$ is the unique weak solution of (2.1) with initial data $u_{0}\in\mathcal{X}$. We note that $\mathcal{X}$ is positively invariant, bounded and closed in $L^{1}(\Omega)$. We can now state our result on the existence of the global attractor for the problem (2.1).
@@
Let (2.2)–(2.9) be satisfied, $\phi(u_{0})\in L^{\infty}(\Omega)$ and $f$ be bounded. Then, the semigroup $\{\mathcal{S}(t)\}_{t\geq 0}$ generated by problem (2.1) has a global attractor $\mathcal{A}$ in $\mathcal{X}$ endowed with the $L^{1}(\Omega)$-topology, which has the form $\mathcal{A}=\omega(\mathcal{X})$.
@@
$\omega(D)=\bigcap_{t>0}\overline{\bigcup_{\tau\geq t}S(t)D}$
represents the $\omega$-limit set of a set $D$ for a semigroup $\{S(t)\}_{t\geq 0}$ in a Banach space $E$, where the bar denotes the closure in $E$.
@@
$\mathcal{Y}:=\overline{\{(u,v)\in L^{1}(\Omega)\times L^{1}(\Omega):\phi(u)\in L^{\infty}(\Omega),v\in[0,1]\}},$
where the closure is taken in $L^{1}(\Omega)\times L^{1}(\Omega)$. Then, by Theorem 2.2, the coupled problem (2.16) generates a semigroup $\{\mathcal{T}(t)\}_{t\geq 0}$ in $\mathcal{Y}$ that is Lipschitz continuous in $L^{1}(\Omega)\times L^{1}(\Omega)$. More specifically, $\mathcal{T}(t)(u_{0},v_{0})=(u(t),v(t))$, where $(u,v)\in C([0,T];L^{1}(\Omega))\times C([0,T];L^{1}(\Omega))$ is the unique weak solution of (2.16) with initial data $(u_{0},v_{0})\in\mathcal{Y}$. Note that $\mathcal{Y}$ is positively invariant, bounded and closed in $L^{1}(\Omega)\times L^{1}(\Omega)$. We now state our result on the existence of the global attractor for the problem (2.16).
@@
Assume that $\phi:[0,1)\to[0,\infty)$ satisfies (2.2)–(2.6) and $f,g$ satisfy (2.17)-(2.18). Moreover, $u^{D}:\Omega\to[0,1)$ satisfies (2.8), $v^{D}:\Omega\to[0,1]$ satisfies $v^{D}\in H^{1}(\Omega)$, and the initial data $u_{0}:\Omega\to[0,1)$ satisfies (2.9), $\phi(u_{0})\in L^{\infty}(\Omega)$ and $v_{0}:\Omega\to[0,1]$ is measurable. Then, the semigroup $\{\mathcal{T}(t)\}_{t\geq 0}$ generated by problem (2.16) has a global attractor $\tilde{\mathcal{A}}$ in $\mathcal{Y}$ endowed with the $L^{1}(\Omega)\times L^{1}(\Omega)$-topology, which has the form $\tilde{\mathcal{A}}=\omega(\mathcal{Y})$.
@@
Theorem 2.4 generalizes the existence result in [12] for the global attractor for the biofilm model (1.3). In fact, we consider the significantly broader class of systems (2.16) and allow for mixed Dirichlet-Neumann boundary data and more general domains. These generalizations are relevant in applications.
@@
For simplicity we assume homogeneous Dirichlet boundary conditions, i.e. $\Gamma=\partial\Omega$ and $u^{D}\equiv 0$ in (2.1). Moreover, the solutions are globally Hölder continuous with exponent $\alpha>0$, i.e. for any solution $u$ of the problem (2.1) there exists a positive constant $M$ such that
$\lVert u\rVert_{C^{\alpha}([0,\infty)\times\overline{\Omega})}\leq M.$ (2.20)
@@
for some constants $C_{1},C_{2}>0$, $\varepsilon\in(0,1)$ and $p>1$ (see [20, Theorem 1.2]). Note that weak solutions $u$ are bounded away from $1$ if the initial data $u_{0}$ has this property. It is expected that these solutions satisfy the global Hölder continuity assumption (2.20), i.e. they are Hölder continuous up to the boundary. In fact, as pointed out by the author in [22, Section 1.7], the proof should extend by using the methods in [6, Chapter III, §11 and §12] if the boundary of the domain, the boundary data and initial data are Hölder continuous.
@@
It was introduced by Kolmogorov and Tihomirov in [23]. The order of growth of $\mathcal{H}_{\varepsilon}^{E}(M)$ as $\varepsilon$ tends to zero is a measure for the massiveness of the set $M$ in $E$. For certain function spaces precise estimates are know. For instance, if $\Omega=(0,1)^{n}$ and $\alpha\in(0,1)$ then
@@
see [5, 9, 11]. In the proof of Theorem 2.5, we use the slightly different construction of the continuous time exponential attractor in [31] and obtain the improved bound
@@
$||(u,v)||_{C^{\tilde{\alpha}}([0,\infty)\times\overline{\Omega})\times C^{\tilde{\alpha}}([0,\infty)\times\overline{\Omega})}\leq\tilde{M}.$ (2.23)
@@
$\text{dim}_{f}^{L^{1}(\Omega)\times L^{1}(\Omega)}(\mathcal{M})\leq\frac{1}{\tilde{\alpha}}+\log_{1/\chi_{1}}\left(N^{W}_{\chi_{2}}\left(B_{1}^{V}(0,0)\right)\right),$
@@ 3 Existence of Global Attractors
In this section, we prove the existence of global attractors for the semigroups generated by problems (2.1) and (2.16), respectively, i.e. we prove Theorems 2.3-2.4. First, we consider the scalar problem (2.1) and show the existence of the global attractor for the semigroup $\{\mathcal{S}(t)\}_{t\geq 0}$ in the phase space $\mathcal{X}\subset L^{1}(\Omega)$, as defined in Section 2. Since $\mathcal{X}$ is positively invariant and bounded, it is certainly a bounded absorbing set for the semigroup. Hence, it suffices to show the asymptotic compactness of the semigroup. To this end, we prove the following lemma by adapting the arguments in [17, §3.5] to our case.

### v2

@@ 2.2 Long-time Behaviour of Solutions
$\mathcal{X}:=\overline{\{u\in L^{1}(\Omega):\|u\|_{L^{\infty}(\Omega)}<1,\phi(u)\in V\}},$
where the bar denotes the closure taken in $L^{1}(\Omega)$. Then, using Theorem 2.1 we will show that problem (2.1) generates a semigroup $\{\mathcal{S}(t)\}_{t\geq 0}$ in $\mathcal{X}$ that is Lipschitz continuous in $L^{1}(\Omega)$. More specifically, $\mathcal{S}(t)u_{0}=u(t)$, where $u\in C([0,T];L^{1}(\Omega))$ is the unique weak solution of (2.1) with initial data $u_{0}\in\mathcal{X}$. We note that $\mathcal{X}$ is positively invariant and bounded in $L^{1}(\Omega)$.
We can now state our result on the existence of the global attractor for the problem (2.1). In addition to (2.2)-(2.6), we assume that $\phi$ satisfies
$|\phi(z)|\leq c_{1}(1-|z|)^{1-a}+c_{2}\qquad\forall z\in I,$ (2.15)
for some constants $a>1$ and $c_{1},c_{2}\geq 0$.
@@
Let (2.2)–(2.8) and (2.15) be satisfied. Then, the semigroup $\{\mathcal{S}(t)\}_{t\geq 0}$ generated by problem (2.1) has a global attractor $\mathcal{A}$ in $\mathcal{X}$ endowed with the $L^{1}(\Omega)$-topology, which has the form $\mathcal{A}=\omega(\mathcal{X})$.
@@
$\omega(D)=\bigcap_{\tau\geq 0}\overline{\bigcup_{t\geq\tau}S(t)D}$
denotes the $\omega$-limit set of a set $D$ for a semigroup $\{S(t)\}_{t\geq 0}$ in a Banach space $E$, where the bar denotes the closure in $E$.
@@
$\mathcal{Y}:=\overline{\{(u,v)\in L^{1}(\Omega)\times L^{1}(\Omega):\|u\|_{L^{\infty}(\Omega)}<1,\phi(u)\in V,u\geq 0,v\in[0,1]\text{ measurable}\}},$
where the bar denotes the closure taken in $L^{1}(\Omega)\times L^{1}(\Omega)$. Then, using Theorem 2.2 we will show that the coupled problem (2.12) generates a semigroup $\{\mathcal{T}(t)\}_{t\geq 0}$ in $\mathcal{Y}$ that is Lipschitz continuous in $L^{1}(\Omega)\times L^{1}(\Omega)$. More specifically, $\mathcal{T}(t)(u_{0},v_{0})=(u(t),v(t))$, where $(u,v)\in C([0,T];L^{1}(\Omega))\times C([0,T];L^{1}(\Omega))$ is the unique weak solution of (2.12) with initial data $(u_{0},v_{0})\in\mathcal{Y}$. Note that $\mathcal{Y}$ is positively invariant and bounded in $L^{1}(\Omega)\times L^{1}(\Omega)$.
We now state our result on the existence of the global attractor for the problem (2.12). As for the scalar problem, we additionally assume that $\phi:[0,1)\to[0,\infty)$ satisfies the inequality in (2.15).
@@
Assume that $\phi:[0,1)\to[0,\infty)$ satisfies (2.2)–(2.6) and (2.15), and $f,g$ satisfy (2.13)-(2.14). Then, the semigroup $\{\mathcal{T}(t)\}_{t\geq 0}$ generated by problem (2.12) has a global attractor $\tilde{\mathcal{A}}$ in $\mathcal{Y}$ endowed with the $L^{1}(\Omega)\times L^{1}(\Omega)$-topology, which has the form $\tilde{\mathcal{A}}=\omega(\mathcal{Y})$.
@@
Theorem 2.4 generalizes the existence result in [21] for the global attractor for the biofilm model (1.3). In fact, we consider the significantly broader class of systems (2.12).
@@
We assume that the solutions are globally Hölder continuous with exponent $\alpha>0$, i.e. for any solution $u$ of the problem (2.1) and compact interval $\widetilde{I}\subset(0,\infty)$ there exists a positive constant $M$ such that
$\lVert u\rVert_{C^{\alpha}(\widetilde{I}\times\overline{\Omega})}\leq M.$ (2.17)
@@
for some constants $C_{1},C_{2}>0$, $\varepsilon\in(0,1)$ and $p>1$ (see [23, Theorem 1.2]). Note that weak solutions $u$ are bounded away from $1$ if the initial data $u_{0}$ has this property. It is expected that these solutions satisfy the global Hölder continuity assumption (2.17), i.e. they are Hölder continuous up to the boundary. In fact, as pointed out by the author in [31, Section 1.7], the proof should extend by using the methods in [32, Chapter III, §11 and §12] if the boundary of the domain, the boundary data and initial data are Hölder continuous.
@@
It was introduced by Kolmogorov and Tihomirov in [33]. The order of growth of $\mathcal{H}_{\varepsilon}^{E}(M)$ as $\varepsilon$ tends to zero is a measure for the massiveness of the set $M$ in $E$. For certain function spaces, precise estimates are known. For instance, if $\Omega=(0,1)^{n}$ and $\alpha\in(0,1)$ then
@@
see [18, 15, 19]. In the proof of Theorem 2.5, we use the slightly different construction of the continuous time exponential attractor in [34], see also [20], and obtain the improved bound
@@
We assume that the solutions are globally Hölder continuous with exponent $\tilde{\alpha}>0$, i.e. for any solution $(u,v)$ of (2.12) and compact interval $\widetilde{I}\subset(0,\infty)$ there exists a positive constant $\tilde{M}$ such that
@@
$\text{dim}_{f}^{L^{1}(\Omega)\times L^{1}(\Omega)}(\mathcal{M})\leq\frac{1}{\tilde{\alpha}}+\log_{1/\chi_{1}}\left(N^{W}_{\chi_{2}}\left(B_{1}^{V}(0,0)\right)\right),$ (2.21)
@@ 3 Existence of Global Attractors
In this section, we prove the existence of global attractors for the semigroups generated by problems (2.1) and (2.12), respectively, i.e. we prove Theorems 2.3 and 2.4. First, we consider the scalar problem (2.1) and show the existence of the global attractor for the semigroup $\{\mathcal{S}(t)\}_{t\geq 0}$ in the phase space $\mathcal{X}\subset L^{1}(\Omega)$ defined in Section 2. This requires uniform bounds for $\phi(u)$, where $u$ is the weak solution in Theorem 2.1. To rigorously justify these estimates, we consider smooth non-degenerate approximations and derive uniform estimates for the approximate solutions exploiting arguments in [1, 31, 21]. Then, by passing to the limit in the approximation parameter, we obtain suitable bounds and smoothing estimates for the weak solution of the original problem and can define the semigroup in $\mathcal{X}$. Since $\mathcal{X}$ is positively invariant and bounded, it is certainly a bounded absorbing set for the semigroup. Hence, it suffices to show the asymptotic compactness of the semigroup which we prove using the uniform estimates for the approximate solutions and by adapting arguments in [24, §3.5] to our case. The results are then extended for the coupled problem.
We first assume that the initial data $u_{0}\in L^{1}(\Omega)$ satisfies
$\|u_{0}\|_{L^{\infty}(\Omega)}\leq 1-\eta,\quad\phi(u_{0})\in V,$ (3.1)
for some $\eta\in(0,1)$, and then generalize the results for initial data $u_{0}\in\mathcal{X}$. We consider the following non-degenerate auxiliary problem
$\begin{cases}\partial_{t}u_{R}=\Delta\phi_{R}(u_{R})+f_{R}(\cdot,u_{R}),&\text{in}\ (0,T]\times\Omega,\\ u_{R}=u_{0,R}&\text{in}\ \{0\}\times\Omega,\\ \phi(u_{R})=0&\text{on}\ (0,T]\times\partial\Omega,\end{cases}$ (3.2)
where $\phi_{R},f_{R}$ and $u_{0,R}$ are smooth approximations for $\phi,f$ and $u_{0}$ with the following properties:
  • $(A_{1})$
    The functions $\phi_{R}\in C^{\infty}(\mathbb{R})$ are convex on $[0,1)$ and concave on (1,0](-1,0], $\phi_{R}\to\phi$ uniformly on compact subsets of $I$, there exist constants $m_{R},M_{R}>0$ such that
    $m_{R}\leq\phi_{R}^{\prime}\leq M_{R},$
    and $\phi_{R}$ is Lipschitz continuous on compact subsets of $I$, where the Lipschitz constant is independent of $R$. Moreover, $|\phi_{R}(z)|\leq|\phi(z)|+\frac{1}{R}|z|$ for any $z\in[-1+1/R,1-1/R]$.
  • $(A_{2})$
    The functions $f_{R}\in C^{\infty}(\overline{\Omega}\times I)$ are Lipschitz continuous,
    $\sup_{z\in I}\|f_{R}(\cdot,z)-f(\cdot,z)\|_{L^{2}(\Omega)}\to 0$
    and $\|f_{R}\|_{L^{\infty}(\Omega\times I)}\leq\|f\|_{L^{\infty}(\Omega\times I)}.$
  • $(A_{3})$
    The functions $u_{0,R}\in C^{\infty}_{c}(\Omega)$ are such that $u_{0,R}\to u_{0}$ in $H^{1}(\Omega)$ and $\|u_{0,R}\|_{L^{\infty}(\Omega)}\leq\|u_{0}\|_{L^{\infty}(\Omega)}\leq 1-\eta.$
The smooth approximations for $\phi$ can be obtained by suitably cutting of the singularity and degeneracy and then taking the convolution with a standard mollifier. Similarly, the smooth approximations for $f$ and $u_{0}$ are constructed.